OpenStudy (anonymous):

Convert the following. y=-2-3 from slope intercept form to standard form

OpenStudy (anonymous):

Is this right -2x-y=-2x-3 -2x-y=-3

OpenStudy (anonymous):

Thank you

OpenStudy (cwrw238):

hold on is the original equation y = -2x - 3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So is it wrong

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

What do I do to fix it

OpenStudy (cwrw238):

add 2x to both sides y + 2x = -2x + 2x - 3 y + 2x = -3

OpenStudy (anonymous):

Isn't the standard form has the x first and not the y

OpenStudy (cwrw238):

oh i guess the x comes before the y :- 2x + y = -3

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

So what would the whole equation be

OpenStudy (cwrw238):

thats it

OpenStudy (anonymous):

So it would be -2x- y=-2x-3

jimthompson5910 (jim_thompson5910):